Define packing fraction and find its value for a body-centred cubic structure


1
Expert's answer
2021-01-27T07:37:33-0500

Solution:


Packing faction or Packing efficiency is the percentage of total space filled by the particles:



In body centred cubic unit cell

In 

Let DF= b

and we know that

ED=EF= a (edge length)

Now,

b2 = a2 + a2 = 2a2

In 

Let, AF = c

We know that

FD = b

& AD = a (edge length)

Now,

c2 = a2 + b2 = a2 + 2a2 = 2a2

or c = "\sqrt{\smash[b]{3}}a"

we know that c is body diagonal. As the sphere at the centre touches the sphere at the corner. Therefore body diagonal c = 4r

i.e. "\sqrt{\smash[b]{3}}a" = 4r

or r = "(\dfrac{\sqrt{\smash[b]{3}}}{4})" a

or a = "\tfrac{4r}{\sqrt{\smash[b]{3}}}"

∴ Volume of the unit cell = a3 = ("\tfrac{4r}{\sqrt{\smash[b]{3}}}")3 = "\tfrac{64r^3}{3\sqrt{\smash[b]{3}}}"

No. of spheres in bcc = 2 

∴ volume of 2 spheres = 2 × "\tfrac{4}{3\pi r^3}"


"Packing \;efficiency=\dfrac{Volume \;occupied \;by\;two\;spheres\;in\;the\;unit\;cell}{Total\;volume\;of\;the\;unit\;cell }*100%"


The packing efficiency of BCC is 68%.